OpenStudy (anonymous):

Jeremy is looking at two storage boxes. The smaller box is a cube with side length s. The length of the larger box, l, is 4 times s. The width of the larger box, w, is 3 times s. The height of the larger box, h, is equal to s. The volume of the larger box, L(h), with height, h, is given with the following function. L(h) = 12h3 Which statement best describes the coefficient of h3? the number of small boxes that will fit along the length of the large box the number of small boxes that will fit inside the large box the area of the base of the large box the increase

OpenStudy (wolf1728):

Volume of Large Box = 4*s * 3*s * s Volume = 12*s^3 So 12 * h3 = 12 * s^3 h3 = s^3 Therefore, h3 is the volume of the smaller box and it is not one of the choices.

OpenStudy (anonymous):

thnx @wolf1728